OpenStudy (anonymous):

solve for x i need help -4(3x-2)=6x+2

OpenStudy (anonymous):

-4(3x-2)=6x+2 -12x+8=6x+2 +12x +12x 8=18x+2 -2 -2 6=18x x=6/18 ---->x=1/3